# Anomalous dimensions from conformal field theory: generalized $\phi^{2n+1}$ theories

### v2

@@ Appendix H Generalized self-avoiding walks, loop-erased random walks and their $\phi^{2n}$ generalizations
$\Delta_{{}_{\mathcal{J}_{\text{S}}}}\big{|}_{N\rightarrow 0}=\;$ $\Delta_{{}_{\mathcal{J}_{\text{ST}}}}\big{|}_{N\rightarrow 0}$
=\;= $d-2k+2m+\ell+\frac{((n-1)!)^{2}(n-1)^{2}\left(1+\frac{k}{n-1}\right)_{k}}{2(n)_{n}^{2}\,{}_{3}F_{2}\left(1-n,\frac{1}{2}-\frac{n}{2},-\frac{n}{2};1,\frac{1}{2}-n;1\right)^{2}}$
×$\times\left(\frac{(-1)^{k+1}(n-1)}{k^{2}\left(1+\frac{nk}{n-1}\right)_{k-1}}-\frac{(k-1)!(2n-1)\left(1+\frac{k}{n-1}\right)_{k}}{(k)_{k-m}(-m)_{m}\left(\ell+m-\frac{(n-2)k}{n-1}\right)_{2k}}\right)+O(\epsilon^{3})\,.$ (184)
@@
  • $\Delta_{{}_{\mathcal{J}_{\text{S}}}}|_{N\rightarrow 0}=\Delta_{{}_{\mathcal{J}_{\text{ST}}}}|_{N\rightarrow 0}$
These correspond to the case of singular tensor structure above. $\Delta_{{}_{\mathcal{J}_{\text{S}}}}|_{N\rightarrow-2}=\Delta_{{}_{\mathcal{J}_{\text{A}}}}|_{N\rightarrow-2}=d-2k+2m+\ell+O(\epsilon^{3})$ Here A denotes the antisymmetric representation. These degeneracies occur due to the vanishing of anomalous dimensions. $\Delta_{{}_{\mathcal{J}_{\text{all three reps}}}}|_{N\rightarrow-4,-6,\ldots,-2n+2}=d-2k+2m+\ell+O(\epsilon^{3})\qquad(n>2)$ These degeneracies are similar to those in the $N\rightarrow-2$ case, except that here all three types of bilinear operators are degenerate. We emphasize that the spin $\ell$ is arbitrary, so all the degeneracies above involve the entire Regge trajectories. The latter two scenarios also lead to $\gamma_{\phi}|_{N\rightarrow-2,-4,\ldots,-2n+2}=O(\epsilon^{3})\,.$ (185) We also comment on the Chew-Frautschi plots. In the $k=1$ case, the positive vertical-axis intercepts associated with $\Delta=d/2$ are ±$\ell^{*}_{\pm}=\frac{n-2}{n-1}+\left(\frac{1}{2}\pm\frac{n\sqrt{r}}{2^{n-3}(n)_{n}\,{}_{3}F_{2}\left(\frac{1}{2}-\frac{n}{2},-\frac{n}{2},-n-\frac{N}{2}+1;1,\frac{1}{2}-n;1\right)}\right)\epsilon+O(\epsilon^{2})\,,$ (186) where the ratio of OPE coefficients in the O($N$)-symmetric $\phi^{2n}$ theory is $r=\begin{cases}4^{n-1}(2n-1)(n-1)!\left(\frac{N}{2}+1\right)_{n-1}&\quad\!\!\!\text{singlet}\\ 2^{2n-3}(4n+N-2)(n-1)!\left(\frac{N}{2}+2\right)_{n-2}&\quad\!\!\!\text{symmetric traceless}\\ 4^{n-1}(n-1)!\left(\frac{N}{2}+1\right)_{n-1}&\quad\!\!\!\text{anti-symmetric.}\end{cases}$ (187) From this perspective, the limit $N\rightarrow-2$ is special in the sense that the two intercepts $\ell^{*}_{\pm}$ become identical for the singlet and antisymmetric representations respectively, and the limits $N\rightarrow-4,\ldots,-2n+2$ are special because $\ell^{*}_{+}=\ell^{*}_{-}$ for all three representations. For $k>1$, there are three qualitative differences. First, the Chew-Frautschi plots contain more curves. Second, there exist trajectories associated with higher $m$. Third, for each $m$, the equation $\Delta=d/2$ has more solutions, indicating that more intersections of the trajectory with the vertical axis can be found. These features are similar to those in the generalized Potts model, which are discussed at the end of subsection 3.1.

### v3

@@ Appendix H Generalized self-avoiding walks, loop-erased random walks, and their $\phi^{2n}$ generalizations
$\Delta_{{}_{\mathcal{J}_{\text{S}}}}\big{|}_{N\rightarrow 0}=\;$ $\Delta_{{}_{\mathcal{J}_{\text{ST}}}}\big{|}_{N\rightarrow 0}=d-2k+2m+\ell+\frac{((n-1)!)^{2}(n-1)^{2}\left(1+\frac{k}{n-1}\right)_{k}}{2(n)_{n}^{2}\,{}_{3}F_{2}\left(1-n,\frac{1}{2}-\frac{n}{2},-\frac{n}{2};1,\frac{1}{2}-n;1\right)^{2}}$
×$\hskip 20.00003pt\times\left(\frac{(-1)^{k+1}(n-1)}{k^{2}\left(1+\frac{nk}{n-1}\right)_{k-1}}-\frac{(-1)^{m}(k-1)!(2n-1)\left(1+\frac{k}{n-1}\right)_{k}}{(k)_{k-m}\,m!\left(\ell+m-\frac{(n-2)k}{n-1}\right)_{2k}}\right)+O(\epsilon^{3})\,.$ (186)
@@
  • $\Delta_{{}_{\mathcal{J}_{\text{S}}}}|_{N\rightarrow 0}=\Delta_{{}_{\mathcal{J}_{\text{ST}}}}|_{N\rightarrow 0}$:
These correspond to the case of the singular tensor structure above. $\Delta_{{}_{\mathcal{J}_{\text{S}}}}|_{N\rightarrow-2}=\Delta_{{}_{\mathcal{J}_{\text{A}}}}|_{N\rightarrow-2}=d-2k+2m+\ell+O(\epsilon^{3})$: Here A denotes the antisymmetric representation. These degeneracies occur due to the vanishing of anomalous dimensions. $\Delta_{{}_{\mathcal{J}_{\text{all three reps}}}}|_{N\rightarrow-4,-6,\ldots,-2n+2}=d-2k+2m+\ell+O(\epsilon^{3}),\quad(n>2)$: These degeneracies are similar to those in the $N\rightarrow-2$ case, except that here all three types of bilinear operators are degenerate. We emphasize that the spin $\ell$ is arbitrary, so all of the degeneracies above involve the entire Regge trajectories. The latter two scenarios also lead to $\gamma_{\phi}|_{N\rightarrow-2,-4,\ldots,-2n+2}=O(\epsilon^{3})\,.$ (187) We also comment on the Chew-Frautschi plots. In the $k=1$ case, the positive vertical-axis intercepts associated with $\Delta=d/2$ are ±$\ell^{*}_{\pm}=\frac{n-2}{n-1}+\left(\frac{1}{2}\pm\frac{n\sqrt{r}}{2^{n-3}(n)_{n}\,{}_{3}F_{2}\left(\frac{1}{2}-\frac{n}{2},-\frac{n}{2},-n-\frac{N}{2}+1;1,\frac{1}{2}-n;1\right)}\right)\epsilon+O(\epsilon^{2})\,,$ (188) where the ratio of OPE coefficients in the O($N$)-symmetric $\phi^{2n}$ theory is $r=\begin{cases}4^{n-1}(2n-1)(n-1)!\left(\frac{N}{2}+1\right)_{n-1}&\quad\!\!\!\text{singlet,}\\ 2^{2n-3}(4n+N-2)(n-1)!\left(\frac{N}{2}+2\right)_{n-2}&\quad\!\!\!\text{symmetric traceless,}\\ 4^{n-1}(n-1)!\left(\frac{N}{2}+1\right)_{n-1}&\quad\!\!\!\text{antisymmetric.}\end{cases}$ (189) From this perspective, the limit $N\rightarrow-2$ is special in the sense that the two intercepts $\ell^{*}_{\pm}$ become identical for the singlet and antisymmetric representations, respectively, and the limits $N\rightarrow-4,\ldots,-2n+2$ are special because $\ell^{*}_{+}=\ell^{*}_{-}$ for all three representations. For $k>1$, there are three qualitative differences. First, the Chew-Frautschi plots contain more curves. Second, there exist trajectories associated with higher $m$. Third, for each $m$, the equation $\Delta=d/2$ has more solutions, indicating that more intersections of the trajectory with the vertical axis can be found. These features are similar to those in the generalized Potts model, which are discussed at the end of section 3.1.